Suppose a sodium atom has a de Broglie wavelength of 1.02 picometer. At what linear speed is it moving, to three significant digits? (The molar mass of Na is 22.990 g/mol.)
1
Expert's answer
2020-07-02T05:41:13-0400

Solution:

1.Find the mass of the sodium atom:

m= M/Na = 22.99/6.02*1023 = 3.819*10-23g = 3.819*10-26kg

2.By the ratio of De Broglie:

α = h/mu; u = h/mα = 6.626*10-34/3.819*10-26*1.02*10-12 = 17010 m/s = 1.70*104m/s

Answer: 1.70*104 m/s